OpenStudy (anonymous):

Evaluate 3t(t+2)-3t^2 for t=19 and Simplifiy -(3a-2b)-3(-a-b) I've already completed them on homework but got them wrong, can someone explain them to me please?

myininaya (myininaya):

Can you show me what you did for the first one? So I can check it please.

OpenStudy (anonymous):

I filled in the t's. So got the equation 3(19)(19+2)-3(19)^2 multiplied 3(19) and 3(19)^2 and got the equation 57(19+2)- 3249 I then used the distributive property gettng 1083+114-3249 and got my final answer -2052. My book says its 114.

myininaya (myininaya):

hmmm... I'm getting 3(19)^2=1083 not 3249

OpenStudy (anonymous):

Okay, yes I got that the second time when I entered it into my calculator just like that.

myininaya (myininaya):

So it should be 57(19+2)-1083 right?

OpenStudy (anonymous):

yes ma'am.

myininaya (myininaya):

Now you were suppose to apply ordered of operations here but for this one it won't end up to matter so far (but you need to be extremely careful with the next steps you take) So Now going back to order of operations It tells us to do what is in the grouping symbols which is 19+2

myininaya (myininaya):

19+2=?

OpenStudy (anonymous):

1197-1083=114!! :) thank you! can you help with the second one? Please & Thank you.

myininaya (myininaya):

lol oh you got it that was quick ok sure

myininaya (myininaya):

-(3a-2b)-3(-a-b) Do you remember the distributive property?

OpenStudy (anonymous):

Yes, and I get -3a--2b-4a-3b.

myininaya (myininaya):

x(r+s)=xr+xs <----> Distributive Property

myininaya (myininaya):

Ok and --=+

myininaya (myininaya):

And also the second pair you multiply a little wrong

OpenStudy (anonymous):

so it should be 2b? not -2b?

myininaya (myininaya):

We have -3(-a-b)